### v1

# Symplectic Generative Networks (SGNs): A Hamiltonian Framework for Invertible Deep Generative Modeling

Abstract
We introduce the Symplectic Generative Network (SGN), a deep generative model that leverages Hamiltonian mechanics to construct an invertible, volume-preserving mapping between a latent space and the data space. By endowing the latent space with a symplectic structure and modeling data generation as the time evolution of a Hamiltonian system, SGN achieves exact likelihood evaluation without incurring the computational overhead of Jacobian determinant calculations. In this work, we provide a rigorous mathematical foundation for SGNs through a comprehensive theoretical framework that includes: (i) complete proofs of invertibility and volume preservation, (ii) a formal complexity analysis with theoretical comparisons to Variational Autoencoders and Normalizing Flows, (iii) strengthened universal approximation results with quantitative error bounds, (iv) an information-theoretic analysis based on the geometry of statistical manifolds, and (v) an extensive stability analysis with adaptive integration guarantees. These contributions highlight the fundamental advantages of SGNs and establish a solid foundation for future empirical investigations and applications to complex, high-dimensional data.
Keywords: Symplectic Generative Networks, Hamiltonian Dynamics, Invertible Neural Networks, Volume-Preserving Flows, Exact Likelihood Estimation, Symplectic Integrators
MSC 2020 Subject Classification: 68T07, 37J39, 65P10, 62B10, 53D22, 94A17
1 Introduction
Deep generative models have revolutionized the field of machine learning by enabling the synthesis and analysis of complex, high-dimensional data. Notably, Variational Autoencoders (VAEs) (Kingma et al. 2014) and Normalizing Flows (NFs) (Rezende et al. 2015) have emerged as powerful frameworks that balance tractable likelihood estimation with expressive latent representations. VAEs achieve efficiency through variational inference, yet their reliance on approximate posteriors can lead to information loss and suboptimal sample quality. In contrast, Normalizing Flows provide exact likelihoods by constructing sequences of invertible transformations; however, the associated computation of Jacobian determinants often results in substantial computational overhead, especially as dimensionality increases.
Despite these successes, a key shortcoming persists in current generative models: many fail to account for the inherent geometric and physical properties of the data. In many real-world scenarios such as those governed by conservation laws in physics, the underlying dynamics exhibit symplectic structure, which guarantees volume preservation and energy conservation. Ignoring these properties may lead to models that are less interpretable and inefficient when applied to structured or physical data.
Motivated by these limitations, we introduce the Symplectic Generative Network (SGN). SGNs reinterpret the latent-to-data mapping as the time evolution of a Hamiltonian (Hamilton 1834) system, thereby naturally embedding the principles of symplectic geometry into the generative process. This approach confers several significant benefits: (i) it guarantees invertibility and exact volume preservation, enabling exact likelihood evaluation without the burdensome computation of Jacobian determinants; (ii) it offers physical interpretability, as the latent dynamics mimic the evolution of systems governed by conservation laws; and (iii) it leverages well-studied numerical integration techniques to ensure stability and efficiency.
In this paper, we provide a rigorous mathematical treatment of SGNs. Our contributions include complete proofs of key properties such as invertibility and volume preservation, formal complexity analyses that compare SGNs to state-of-the-art models like VAEs and Normalizing Flows, and strengthened universal approximation theorems with quantitative error bounds. We further develop an information-theoretic analysis rooted in the geometry of statistical manifolds and present extensive stability guarantees through adaptive integration strategies. Collectively, these results not only deepen the theoretical understanding of SGNs but also lay a solid foundation for future empirical research and applications across diverse domains, from physical systems and time series to structured data.
The remainder of this paper is organized as follows. Section 2 reviews related work. Section 3 defines the SGN model, while Section 4 provides the theoretical analysis of SGN. Section 5 presents a comparison between SGN and existing generative models. Section 6 presents strengthened universal approximation results, and Section 7 provides the information-theoretic analysis. Section 8 gives the expanded stability analysis, and Section 9 provides the pseudocode for the SGN training procedure. Finally, Section 10 concludes with remarks and directions for future research.
2 Related Work
The past decade has witnessed remarkable advances in deep generative modeling, driven by the need to balance expressivity, computational efficiency, and tractable likelihood evaluation. Early frameworks such as Variational Autoencoders (VAEs) (Kingma et al. 2014) laid the groundwork by combining probabilistic inference with deep learning. While VAEs benefit from computational tractability and ease of implementation, their reliance on variational approximations introduces a gap between the true posterior and its approximation, often resulting in information loss and less accurate sample generation.
Normalizing Flows (NFs) (Rezende et al. 2015) were developed to overcome these limitations by constructing invertible transformations that yield exact likelihoods. Despite their theoretical appeal, normalizing flows typically incur high computational costs due to the need to compute Jacobian determinants, costs that can scale poorly with data dimensionality. Recent efforts, such as the development of residual flows (Chen et al. 2019), have aimed to mitigate this issue; however, these methods still face challenges when scaling to very high dimensions or when complex transformations are required.
A more recent line of work has sought to integrate physical principles into deep learning models. Hamiltonian Neural Networks (Greydanus et al. 2019) have demonstrated that incorporating Hamiltonian dynamics can improve the modeling of time-evolving systems by enforcing energy conservation and other physical laws. Building on this idea, SGNs embed a symplectic integrator directly into the generative framework. This approach ensures not only invertibility and exact likelihood evaluation but also a natural adherence to the conservation properties intrinsic to many real-world phenomena.
In addition, reversible architectures (Gomez et al. 2017) have emphasized memory efficiency during backpropagation by leveraging the reversibility of the network, a property that is central to SGNs. By grounding our method in the principles of symplectic geometry, we provide a mathematically rigorous pathway for integrating physical dynamics into deep generative modeling. This synthesis of ideas from physics, numerical analysis, and deep learning opens new avenues for research in generative models that are both theoretically robust and practically efficient.
Overall, our work represents a significant step toward generative models that are not only expressive and computationally efficient but also aligned with the underlying geometric and physical structure of the data. This alignment has the potential to yield models that are more interpretable and better suited for applications in scientific computing, time series analysis, and beyond.
3 Symplectic Generative Networks (SGNs)
In this section, we introduce the Symplectic Generative Network (SGN), a novel deep generative framework that integrates Hamiltonian dynamics into the latent-to-data mapping. SGNs leverage the intrinsic properties of Hamiltonian systems such as volume preservation, invertibility, and energy conservation to achieve exact likelihood evaluation and enhanced interpretability. Our model is structured around three key components: (i) a latent space defined via canonical coordinates, (ii) a symplectic integrator that evolves the latent state according to Hamilton’s equations, and (iii) a variational inference scheme employing a Hamiltonian encoder. We describe these components in detail below.
3.1 Latent Space and Hamiltonian Dynamics
We begin by defining the latent space $\mathcal{Z}$ as
$z\in\mathbb{R}^{2d},$
which is interpreted as the concatenation of canonical coordinates:
$z=(q,p),\quad q,p\in\mathbb{R}^{d}.$
In this formulation, $q$ and $p$ represent generalized positions and momenta, respectively. We assume a standard Gaussian prior on the latent space,
$p(z)=\mathcal{N}(0,I_{2d}),$
providing a simple and well-behaved distribution for sampling.
To model the data generation process, we introduce a neural network $H_{\psi}(q,p)$, parameterized by $\psi$, that serves as the Hamiltonian of the system. The dynamics of the latent state are governed by Hamilton’s equations:
$\dot{q}=\frac{\partial H_{\psi}}{\partial p},\quad\dot{p}=-\frac{\partial H_{\psi}}{\partial q}.$ (1)
For a fixed integration time $T$, the solution to these differential equations defines a flow map $\Phi_{T}$ that transforms an initial latent state $z_{0}=(q_{0},p_{0})$ into a new state $z_{T}=(q_{T},p_{T})$:
$(q_{T},p_{T})=\Phi_{T}(q_{0},p_{0};\psi).$
This formulation guarantees that the latent-to-data transformation is both invertible and volume preserving, thanks to the inherent symplectic structure.
To provide a concrete illustration of Hamiltonian dynamics, we include the Figure 1 which depicts phase space trajectories. The blue circles represent constant energy levels, while the red trajectory shows how a system evolves according to Hamilton’s equations:
$q$$p$Energy LevelsHamiltonian Flow Hamilton’s Equations:
$q^{\prime}=\frac{\partial H}{\partial p},\quad p^{\prime}=-\frac{\partial H}{\partial q}$
Preserves symplectic form.
Figure 1: Hamiltonian Dynamics in Phase Space. Blue circles denote energy levels, while the red trajectory illustrates the evolution under Hamilton’s equations, highlighting volume preservation and energy conservation.
The overall generative process in SGNs is as follows:
  1. 1.
    Sampling: Draw an initial latent state $z_{0}=(q_{0},p_{0})$ from the prior $p(z)$.
  2. 2.
    Evolution: Evolve $z_{0}$ using the Hamiltonian flow $\Phi_{T}$ to obtain $z_{T}=\Phi_{T}(z_{0};\psi)$.
  3. 3.
    Decoding: Generate data $x$ from the decoder $p_{\theta}(x\mid z_{T})$.
By interpreting the latent dynamics through Hamiltonian mechanics, SGNs naturally incorporate physical principles into deep generative modeling.
3.2 Symplectic Integrator
A central component of SGNs is the use of a symplectic integrator to approximate the continuous Hamiltonian flow. We employ the leapfrog (Stoermer–Verlet (Verlet 1967)) integrator due to its favorable properties: it is a second-order method that is both time-reversible and volume preserving. The preservation of the canonical symplectic form,
$\omega=dq\wedge dp,$
ensures that the flow $\Phi_{T}$ satisfies
$\left|\det\frac{\partial\Phi_{T}(z_{0})}{\partial z_{0}}\right|=1,$
eliminating the need for costly Jacobian determinant computations during likelihood evaluation.
The leapfrog integrator advances the state $(q,p)$ through the following three-step update with step size $\Delta t$:
$p_{t+\frac{1}{2}}$ $=p_{t}-\frac{\Delta t}{2}\frac{\partial H_{\psi}}{\partial q}(q_{t},p_{t}),$ (2)
$q_{t+1}$ $=q_{t}+\Delta t\,\frac{\partial H_{\psi}}{\partial p}(q_{t},p_{t+\frac{1}{2}}),$ (3)
$p_{t+1}$ $=p_{t+\frac{1}{2}}-\frac{\Delta t}{2}\frac{\partial H_{\psi}}{\partial q}(q_{t+1},p_{t+\frac{1}{2}}).$ (4)
Iterating these steps $N$ times (with $T=N\Delta t$) produces the discrete flow map $\Phi_{T}$. The second-order accuracy and inherent symplecticity of the leapfrog method ensure that the numerical integration preserves the physical and geometric properties of the latent space.
We now include a schematic diagram in Figure 2 of the overall SGN architecture to provide a high-level overview of the model:
Data Space $x$ Encoder $q_{\phi}(z_{0}\mid x)$ Hamiltonian $H_{\psi}(q,p)$ Decoder $p_{\theta}(x\mid z_{T})$ Prior $\mathcal{N}(0,I)$ $(q_{0},p_{0})$$z_{T}=\Phi_{T}(z_{0})$Reverse Flow
Figure 2: SGN Architecture: The encoder maps data $x$ to an initial latent state $(q_{0},p_{0})$, which is evolved via a Hamiltonian flow $\Phi_{T}$ into $z_{T}$. The decoder then maps $z_{T}$ back to the data space. The symplectic nature of the flow ensures exact volume preservation and invertibility.
Figure 3 shows the leapfrog integration in SGNs.
Step 1: $p_{t+\frac{1}{2}}=p_{t}-\frac{\Delta t}{2}\nabla_{q}H(q_{t},p_{t})$ Step 2: $q_{t+1}=q_{t}+\Delta t\,\nabla_{p}H(q_{t},p_{\frac{1}{2}})$ Step 3: $p_{t+1}=p_{\frac{1}{2}}-\frac{\Delta t}{2}\nabla_{q}H(q_{t+1},p_{\frac{1}{2}})$ Properties:
$\bullet$ Symplectic (volume preserving)
$\bullet$ Reversible and second-order accurate
$\bullet$ Local Error: $O(\Delta t^{3})$, Global Error: $O(T\Delta t^{3})$
Figure 3: Leapfrog Integration in SGNs. The three-step update ensures a symplectic (volume-preserving) evolution of the latent state, enabling exact likelihood evaluation without Jacobian determinants.
3.3 Variational Inference with a Hamiltonian Encoder
To perform inference and learn meaningful latent representations, SGNs incorporate a stochastic encoder $q_{\phi}(z_{0}\mid x)$ that maps observed data $x$ to an initial latent state $z_{0}=(q_{0},p_{0})$. Parameterized by $\phi$, the encoder captures the uncertainty inherent in mapping from data to latent space. Once $z_{0}$ is obtained, it is deterministically evolved via the symplectic flow $\Phi_{T}$ to produce the final latent state $z_{T}$. A decoder network $p_{\theta}(x\mid z_{T})$ then reconstructs or generates data from $z_{T}$.
The overall variational objective is given by the Evidence Lower Bound (ELBO):
$\mathcal{L}_{\mathrm{SGN}}(x)=\mathbb{E}_{q_{\phi}(z_{0}\mid x)}\Bigl{[}\log p_{\theta}\bigl{(}x\mid\Phi_{T}(z_{0};\psi)\bigr{)}\Bigr{]}-D_{\mathrm{KL}}\Bigl{(}q_{\phi}(z_{0}\mid x)\,\|\,p(z_{0})\Bigr{)}.$ (5)
Since the symplectic flow $\Phi_{T}$ is volume preserving (its Jacobian determinant equals 1), no additional correction is needed when transforming from $z_{0}$ to $z_{T}$. This results in an exact likelihood evaluation that is both computationally efficient and theoretically robust.
In summary, the SGN framework integrates a carefully designed latent space, a robust symplectic integrator, and a probabilistic inference scheme to construct a deep generative model that is computationally efficient, theoretically sound, and physically interpretable. The integration of Hamiltonian dynamics not only preserves essential geometric properties but also aligns the model with underlying physical principles, making SGNs particularly attractive for applications in domains governed by conservation laws.
4 Theoretical Analysis
4.1 Invertibility and Volume Preservation
Theorem 4.1 (Symplecticity and Volume Preservation).
Let $\Phi_{T}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ be the flow obtained by integrating (1) using a symplectic integrator with step size $\Delta t$ over $T$ steps. Then, $\Phi_{T}$ is invertible and volume preserving:
$\left|\det\frac{\partial\Phi_{T}(z_{0})}{\partial z_{0}}\right|=1,\quad\forall\,z_{0}\in\mathbb{R}^{2d}.$
Proof.
A mapping $\Phi:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ is symplectic if it preserves the canonical form
$\omega=\sum_{i=1}^{d}dq_{i}\wedge dp_{i}.$
This is equivalent to requiring that
$D\Phi(z)^{T}\,J\,D\Phi(z)=J,$
where
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}.$
Taking determinants gives
$\det\bigl{(}D\Phi(z)^{T}\,J\,D\Phi(z)\bigr{)}=\det(J)=1.$
Since $\det(D\Phi(z)^{T})=\det(D\Phi(z))$, it follows that
$\det(D\Phi(z))^{2}=1\quad\Longrightarrow\quad|\det(D\Phi(z))|=1.$
Moreover, since the leapfrog integrator is constructed from shear maps (each with unit determinant), their composition yields a unit Jacobian. ∎
4.2 Exact Likelihood Evaluation
Because $\Phi_{T}$ is volume preserving, the likelihood becomes:
[…]
Under the change of variables $z_{T}=\Phi_{T}(z_{0})$, the Jacobian term is unity, enabling exact likelihood computation.
4.3 Stability and Expressivity Analysis
The neural network $H_{\psi}(q,p)$ is designed to be highly expressive. Its gradients dictate the latent evolution, and the leapfrog integrator’s local error is $\mathcal{O}(\Delta t^{3})$ (global error $\mathcal{O}(T\Delta t^{3})$). For example, for the quadratic Hamiltonian
$H(q,p)=\frac{1}{2}p^{2}+\frac{\omega^{2}}{2}q^{2},$
stability requires $\Delta t\,\omega<2$. For general $H_{\psi}$, local frequencies may be estimated from the Hessian’s eigenvalues, and adaptive or higher-order methods can improve stability.
5 Theoretical Comparison with Existing Generative Models
5.1 Formal Analysis of Computational Complexity
Theorem 5.1 (Complexity Advantage of SGNs).
Let $d$ be the latent dimensionality and $C_{J}(d)$ the cost for computing a $d\times d$ Jacobian determinant. For a normalizing flow with $K$ coupling layers, the exact log-likelihood evaluation requires $\mathcal{O}(K\cdot C_{J}(d))$ operations, while for an SGN with $T$ integration steps the cost is $\mathcal{O}(T\cdot d)$, independent of any Jacobian computation.
Proof.
In a normalizing flow, the mapping from data $x$ to the latent variable $z_{K}$ is given by a sequence of $K$ invertible transformations:
$z_{K}=f_{K}\circ f_{K-1}\circ\cdots\circ f_{1}(z_{0}),$
where each $f_{i}$ is an invertible transformation (often implemented as a coupling layer). By the change-of-variables formula, the log-likelihood is computed as
$\log p(x)=\log p(z_{K})+\sum_{i=1}^{K}\log\left|\det\frac{\partial f_{i}}{\partial z_{i-1}}\right|,$
where $z_{i-1}=f_{i-1}\circ\cdots\circ f_{1}(z_{0})$.
Assume that computing the determinant of the Jacobian matrix $\frac{\partial f_{i}}{\partial z_{i-1}}$, which is a $d\times d$ matrix, requires $C_{J}(d)$ operations. Since there are $K$ such layers, the total computational cost for these determinant evaluations is
$\mathcal{O}(K\cdot C_{J}(d)).$
Now consider the SGN framework. SGNs use a symplectic integrator (e.g., the leapfrog method) to simulate the Hamiltonian dynamics defined by
$\dot{q}=\frac{\partial H_{\psi}}{\partial p},\quad\dot{p}=-\frac{\partial H_{\psi}}{\partial q},$
where $z=(q,p)$. The integrator discretizes the continuous-time evolution into $T$ steps with step size $\Delta t$. In each integration step, the following updates are performed:
  1. 1.
    Half-step update for $p$:
    $p_{t+\frac{1}{2}}=p_{t}-\frac{\Delta t}{2}\frac{\partial H_{\psi}}{\partial q}(q_{t},p_{t}).$
  2. 2.
    Full-step update for $q$:
    $q_{t+1}=q_{t}+\Delta t\,\frac{\partial H_{\psi}}{\partial p}(q_{t},p_{t+\frac{1}{2}}).$
  3. 3.
    Another half-step update for $p$:
    $p_{t+1}=p_{t+\frac{1}{2}}-\frac{\Delta t}{2}\frac{\partial H_{\psi}}{\partial q}(q_{t+1},p_{t+\frac{1}{2}}).$
Assuming that evaluating a gradient (either $\frac{\partial H_{\psi}}{\partial q}$ or $\frac{\partial H_{\psi}}{\partial p}$) requires $\mathcal{O}(d)$ operations, the cost per integration step is $\mathcal{O}(d)$ (since each step performs a constant number of gradient evaluations). With $T$ integration steps, the total cost is
$\mathcal{O}(T\cdot d).$
Furthermore, due to the symplectic property of the integrator, we have
$\left|\det\frac{\partial\Phi_{T}}{\partial z_{0}}\right|=1,$
which means that there is no need to compute any additional Jacobian determinants.
Thus, we conclude that the overall computational cost for SGNs is $\mathcal{O}(T\cdot d)$, which is independent of any expensive Jacobian computations. Comparing both approaches, the computational advantage of SGNs is established. ∎
Proposition 5.2 (Memory Complexity).
The memory complexity during backpropagation for SGNs is $\mathcal{O}(T+d)$ (by leveraging reversibility), compared to $\mathcal{O}(K\cdot d)$ for normalizing flows.
Proof.
In normalizing flows, the forward pass involves a sequence of $K$ coupling layers. During backpropagation, one must store the activations (or intermediate outputs) from each of these layers to compute gradients, which leads to a memory requirement that scales as $\mathcal{O}(K\cdot d)$, where $d$ denotes the dimensionality of the latent space.
In contrast, SGNs are built using a reversible (symplectic) integrator. The key property of such integrators is that the forward computation is invertible, allowing the reconstruction of intermediate states during the backward pass rather than storing them explicitly. Consequently, one only needs to store the current state and minimal auxiliary information (such as gradients), resulting in a memory complexity of $\mathcal{O}(T+d)$, where $T$ is the number of integration steps and $d$ is the latent dimensionality. This reduction in memory footprint is a significant advantage for SGNs, particularly when $K$ is large. ∎
5.2 Theoretical Bounds on Approximation Capabilities
Theorem 5.3 (Expressivity Comparison).
Let $\mathcal{M}_{2d}$ denote the set of volume-preserving diffeomorphisms on $\mathbb{R}^{2d}$ and $\mathcal{H}_{d}$ the set of Hamiltonian flows. Then:
  1. 1.
    Every $\Phi\in\mathcal{H}_{d}$ preserves volume, i.e., $\mathcal{H}_{d}\subset\mathcal{M}_{2d}$.
  2. 2.
    Not every volume-preserving map is Hamiltonian, i.e., $\mathcal{H}_{d}\subsetneq\mathcal{M}_{2d}$.
  3. 3.
    However, for any $\Phi\in\mathcal{M}_{2d}$ isotopic to the identity, there exists a sequence of Hamiltonian flows that uniformly approximate $\Phi$ on compact sets.
Proof.
(1) Hamiltonian flows preserve volume:
By Liouville’s theorem, any Hamiltonian flow generated by a smooth Hamiltonian $H(q,p)$ preserves the canonical symplectic form
$\omega=\sum_{i=1}^{d}dq_{i}\wedge dp_{i}.$
Preservation of this form implies that the Jacobian determinant of the flow satisfies
$\left|\det\frac{\partial\Phi}{\partial(q,p)}\right|=1,$
which is exactly the condition for volume preservation. Hence, every $\Phi\in\mathcal{H}_{d}$ is also in $\mathcal{M}_{2d}$.
(2) Not every volume-preserving map is Hamiltonian:
Consider a shear mapping defined on $\mathbb{R}^{2}$ by
$S(x,y)=(x+f(y),\,y),$
where $f$ is a smooth function. This map has a Jacobian determinant of
$\det\begin{pmatrix}1&f^{\prime}(y)\\ 0&1\end{pmatrix}=1,$
so it is volume preserving. However, for $S$ to be Hamiltonian (i.e., generated by some Hamiltonian $H(q,p)$ via Hamilton’s equations), the transformation must preserve the canonical two-form $dq\wedge dp$ in a manner consistent with a Hamiltonian vector field. In general, unless $f$ is linear (which would yield a linear, hence symplectic, transformation), the shear $S$ does not arise from a Hamiltonian flow. Therefore, there exist volume-preserving maps in $\mathcal{M}_{2d}$ that are not Hamiltonian, i.e., $\mathcal{H}_{d}\subsetneq\mathcal{M}_{2d}$.
(3) Uniform approximation by Hamiltonian flows:
Let $\Phi\in\mathcal{M}_{2d}$ be a volume-preserving diffeomorphism isotopic to the identity. By Moser’s theorem, there exists a smooth one-parameter family $\{\Phi_{t}\}_{t\in[0,1]}$ of volume-preserving diffeomorphisms with $\Phi_{0}=\mathrm{Id}$ and $\Phi_{1}=\Phi$. The time-dependent vector field generating this isotopy can be decomposed (via, for instance, the Hodge decomposition) into a Hamiltonian part and a divergence-free remainder. Under suitable conditions, the divergence-free remainder can be made arbitrarily small, so that the vector field is approximated by a Hamiltonian vector field $J\nabla H_{t}$. By employing universal approximation results for neural networks, one can approximate the required Hamiltonian function $H_{t}$ arbitrarily well on any compact set. Consequently, the flow generated by this approximate Hamiltonian (using a symplectic integrator) uniformly approximates $\Phi$. This shows that every volume-preserving diffeomorphism isotopic to the identity can be approximated arbitrarily well by a sequence of Hamiltonian flows. ∎
Proposition 5.4 (Approximation Rate Comparison).
Assume a target diffeomorphism $\Phi$ is approximated by either a normalizing flow with $K$ layers or an SGN with $T$ integration steps and a neural network Hamiltonian of width $n$. Then:
  • Normalizing flows: $\varepsilon_{NF}=\mathcal{O}\left(K^{-1/2}\cdot n^{-1/2}\right)$.
  • SGNs (for volume-preserving targets): $\varepsilon_{SGN}=\mathcal{O}\left(T^{-1}\cdot n^{-1/(2d)}\right)$.
Proof.
We consider the two cases separately.
Normalizing Flows:
Assume that each coupling layer in a normalizing flow approximates a partial transformation with an approximation error of
$\mathcal{O}\left(n^{-1/2}\right)$
in a suitable norm, as suggested by standard universal approximation results for neural networks with width $n$. When $K$ such layers are composed to approximate the target diffeomorphism $\Phi$, the overall error does not simply add up linearly; under reasonable assumptions (e.g., statistical independence or mild interactions between the layers), the errors can accumulate in a root-mean-square fashion. Hence, the total error becomes
$\varepsilon_{NF}=\mathcal{O}\left(\sqrt{\frac{1}{K}}\cdot n^{-1/2}\right)=\mathcal{O}\left(K^{-1/2}\cdot n^{-1/2}\right).$
SGNs:
In SGNs, there are two principal sources of error when approximating a target volume-preserving diffeomorphism:
  1. 1.
    Approximation error of the neural network Hamiltonian: Let $H_{\psi}$ be the neural network approximating the true Hamiltonian underlying $\Phi$. Standard approximation results indicate that for functions defined on $\mathbb{R}^{2d}$ (since $z=(q,p)\in\mathbb{R}^{2d}$), the error in the $C^{1}$ norm decreases as
    $\mathcal{O}\left(n^{-1/(2d)}\right),$
    where $n$ is the network width.
  2. 2.
    Discretization error of the symplectic integrator: The continuous Hamiltonian flow is approximated using a symplectic (e.g., leapfrog) integrator, which introduces a local error of $\mathcal{O}(\Delta t^{3})$ per integration step. Over $T$ steps, with a fixed total integration time, the global integration error scales as
    $\mathcal{O}(T\cdot\Delta t^{3}).$
    By choosing the integration step size $\Delta t$ appropriately (so that $T\Delta t$ is constant), this global error can be balanced with the neural network approximation error. For simplicity, if we assume the integration error is controlled and scales inversely with the number of steps (i.e., $\Delta t\propto T^{-1}$), then the overall discretization error is of order
    $\mathcal{O}\left(T^{-1}\right).$
Combining the two sources, the overall approximation error for SGNs becomes
$\varepsilon_{SGN}=\mathcal{O}\left(T^{-1}\cdot n^{-1/(2d)}\right).$
Thus, we have shown that the approximation errors scale as stated:
$\varepsilon_{NF}=\mathcal{O}\left(K^{-1/2}\cdot n^{-1/2}\right)\quad\text{and}\quad\varepsilon_{SGN}=\mathcal{O}\left(T^{-1}\cdot n^{-1/(2d)}\right).$
Theorem 5.5 (Information Preservation).
Let $X$ be a random variable with distribution $p_{X}$ and let $Z$ denote its latent representation obtained via an invertible mapping $f$ (as in normalizing flows or SGNs). Then:
  1. 1.
    For a deterministic, invertible model, it holds that
    $I(X;Z)=H(X)=H(Z).$
  2. 2.
    For a stochastic model (e.g., VAEs), we have
    $I(X;Z)<H(X).$
Proof.
For an invertible mapping $f\colon\mathcal{X}\to\mathcal{Z}$ where $z=f(x)$, the change-of-variables formula for differential entropy gives
$H(Z)=H(X)+\mathbb{E}_{X}\left[\log\left|\det\frac{\partial f(x)}{\partial x}\right|\right].$
In models such as normalizing flows or SGNs, the mapping $f$ is designed to be volume preserving, meaning that
$\left|\det\frac{\partial f(x)}{\partial x}\right|=1\quad\text{for all }x.$
Thus, the expectation term vanishes:
$\mathbb{E}_{X}\left[\log\left|\det\frac{\partial f(x)}{\partial x}\right|\right]=0,$
and we obtain
$H(Z)=H(X).$
Since $f$ is deterministic and invertible, knowing $Z$ exactly determines $X$ (and vice versa). Therefore, the conditional entropy $H(X\mid Z)=0$, and by the definition of mutual information,
$I(X;Z)=H(X)-H(X\mid Z)=H(X).$
Equivalently, we have
$I(X;Z)=H(Z),$
which establishes the first claim.
In contrast, for stochastic models such as VAEs, the encoder $q_{\phi}(z|x)$ maps an input $x$ to a distribution over latent variables $z$ rather than to a unique $z$. This stochasticity means that the conditional entropy $H(X\mid Z)$ is strictly positive, reflecting the uncertainty in reconstructing $x$ from $z$. As a result, the mutual information satisfies
$I(X;Z)=H(X)-H(X\mid Z)<H(X).$
This demonstrates that stochastic models lose some information during the encoding process. ∎
Corollary 5.6 (Volume Preservation Constraint).
An invertible generative model preserves information if and only if it preserves volume (i.e., has unit Jacobian) or explicitly accounts for volume changes.
Proof.
Let $f$ be an invertible transformation mapping data $x$ to latent representation $z$, i.e., $z=f(x)$. By the change-of-variables formula for differential entropy, we have
$H(Z)=H(X)+\mathbb{E}_{X}\left[\log\left|\det\frac{\partial f(x)}{\partial x}\right|\right].$
If the mapping $f$ preserves volume, then
$\left|\det\frac{\partial f(x)}{\partial x}\right|=1\quad\text{for all }x,$
and thus
$\mathbb{E}_{X}\left[\log\left|\det\frac{\partial f(x)}{\partial x}\right|\right]=0,$
which implies that
$H(Z)=H(X).$
Since the model is invertible, no information is lost and the mutual information satisfies
$I(X;Z)=H(X)-H(X\mid Z)=H(X),$
given that $H(X\mid Z)=0$.
Conversely, if the mapping $f$ does not preserve volume (i.e., the Jacobian determinant is not uniformly one), then the term
$\mathbb{E}_{X}\left[\log\left|\det\frac{\partial f(x)}{\partial x}\right|\right]$
will be non-zero, which results in either an increase or decrease in the differential entropy $H(Z)$ relative to $H(X)$. In such cases, unless the model explicitly corrects for these changes (for example, by incorporating the Jacobian determinant into its likelihood computation), the information in $X$ is not perfectly preserved in $Z$.
Therefore, an invertible generative model preserves information if and only if it either preserves volume (i.e., has unit Jacobian everywhere) or it explicitly accounts for volume changes. ∎
Proposition 5.7 (Trade-off Characterization).
Generative models trade off as follows:
  1. 1.
    VAEs: Low computational complexity but with information loss.
  2. 2.
    Normalizing Flows: High expressivity and exact likelihood but high computational cost.
  3. 3.
    SGNs: Exact likelihood and low computational complexity, with expressivity constrained to volume-preserving maps.
Proof.
VAEs employ approximate inference, typically optimizing a variational lower bound, which leads to an approximate posterior and consequently some loss of information about the data distribution. This results in a lower computational burden but with a trade-off in the fidelity of the representation.
Normalizing flows construct a sequence of invertible transformations that allow for exact likelihood computation. However, to maintain invertibility, these models often require the computation of Jacobian determinants or related quantities, which can be computationally expensive (scaling poorly with the latent dimension in the worst case).
SGNs, by contrast, leverage symplectic integrators to simulate Hamiltonian dynamics. These integrators are designed to be volume preserving (i.e., they have a unit Jacobian), which means that exact likelihood evaluation is achieved without incurring the computational cost of Jacobian determinant calculations. The drawback is that the class of transformations that SGNs can represent is restricted to those that preserve volume. In other words, while SGNs enjoy lower computational cost and exact likelihoods, they are less expressive than normalizing flows when it comes to representing general invertible maps. ∎
6 Strengthened Universal Approximation Results
6.1 Universal Approximation of Volume-Preserving Maps
Theorem 6.1 (Universal Approximation of Volume-Preserving Maps).
Let $\Omega\subset\mathbb{R}^{2d}$ be a compact set and $\Phi:\Omega\to\mathbb{R}^{2d}$ be a $C^{1}$-smooth volume-preserving diffeomorphism isotopic to the identity. Then, for any $\epsilon>0$, there exist:
  1. 1.
    A neural network Hamiltonian $H_{\psi}:\mathbb{R}^{2d}\to\mathbb{R}$ with Lipschitz continuous gradients,
  2. 2.
    A time $T>0$ and step size $\Delta t>0$ with $N=T/\Delta t$,
such that the symplectic flow $\Phi_{T}$ induced by the leapfrog integrator satisfies
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi(z)\|<\epsilon.$
Proof.
We prove the theorem in several steps.
Step 1: Representing the Target Map as a Flow.
Since $\Phi$ is a $C^{1}$ volume-preserving diffeomorphism on the compact set $\Omega$ and is isotopic to the identity, Moser’s theorem guarantees the existence of a smooth one-parameter family of volume-preserving diffeomorphisms $\{\Phi_{t}\}_{t\in[0,1]}$ such that
$\Phi_{0}=\mathrm{Id}\quad\text{and}\quad\Phi_{1}=\Phi.$
Moreover, there exists a time-dependent divergence-free vector field $v_{t}(z)$ on $\Omega$ satisfying
$\frac{d}{dt}\Phi_{t}(z)=v_{t}(\Phi_{t}(z)).$
Because $\Omega$ is a subset of $\mathbb{R}^{2d}$ (which is contractible), every divergence-free vector field is exact. Hence, for each $t$, there exists a smooth function $H_{t}(z)$ such that
$v_{t}(z)=J\nabla H_{t}(z),$
where
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$
is the canonical symplectic matrix. In other words, the target map $\Phi$ can be viewed as the time-$1$ flow of a (possibly time-dependent) Hamiltonian vector field.
Step 2: Approximating the Hamiltonian with a Neural Network.
For a fixed time $t$, by the universal approximation theorem for neural networks, for any $\epsilon_{1}>0$ there exists a neural network $H_{\psi}:\mathbb{R}^{2d}\to\mathbb{R}$ with Lipschitz continuous gradients that approximates the true Hamiltonian $H_{t}$ (or an appropriate average over $t$) uniformly in the $C^{1}$ norm on $\Omega$. That is,
$\sup_{z\in\Omega}\|\nabla H_{\psi}(z)-\nabla H_{t}(z)\|<\epsilon_{1}.$
Consequently, the Hamiltonian vector field $J\nabla H_{\psi}(z)$ approximates $J\nabla H_{t}(z)$ uniformly on $\Omega$.
Step 3: Discretizing the Flow via a Symplectic Integrator.
Let $\Phi^{H_{\psi}}_{t}$ denote the continuous flow generated by the Hamiltonian vector field $J\nabla H_{\psi}(z)$. We now discretize this flow using a symplectic integrator (e.g., the leapfrog method). For a chosen step size $\Delta t$ and total integration time $T$ (with $N=T/\Delta t$), let $\Phi_{T}$ be the discrete flow obtained by iterating the integrator. Standard error analysis for symplectic integrators shows that, for sufficiently small $\Delta t$, there exists $\epsilon_{2}>0$ such that
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi^{H_{\psi}}_{T}(z)\|<\epsilon_{2}.$
Here, $\epsilon_{2}$ can be made arbitrarily small by choosing $\Delta t$ appropriately.
Step 4: Combining the Approximations.
Denote by $\Phi(z)$ the target map, which equals $\Phi_{1}(z)$. Using the triangle inequality, we have
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi(z)\|\leq\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi^{H_{\psi}}_{T}(z)\|+\sup_{z\in\Omega}\|\Phi^{H_{\psi}}_{T}(z)-\Phi(z)\|.$
The first term is bounded by $\epsilon_{2}$ as shown above. The second term reflects the error due to approximating the true Hamiltonian $H_{t}$ by the neural network $H_{\psi}$; by our choice of $\epsilon_{1}$ (and appropriate control over the integration time $T$), this term can be bounded by $\epsilon_{1}$. Therefore, by choosing $\epsilon_{1}$ and $\epsilon_{2}$ such that
$\epsilon_{1}+\epsilon_{2}<\epsilon,$
we obtain
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi(z)\|<\epsilon.$
Conclusion: By the above steps, we have constructed a neural network Hamiltonian $H_{\psi}$ with Lipschitz continuous gradients, and by choosing appropriate integration parameters $T$ and $\Delta t$ (with $N=T/\Delta t$), the symplectic flow $\Phi_{T}$ approximates the target volume-preserving diffeomorphism $\Phi$ uniformly on $\Omega$ to within any pre-specified error $\epsilon>0$. ∎
6.2 Quantitative Bounds on Approximation Error
Theorem 6.2 (Quantitative Approximation Bounds).
Let $\Phi:\Omega\to\mathbb{R}^{2d}$ be a $C^{2}$-smooth volume-preserving diffeomorphism on a compact set $\Omega\subset\mathbb{R}^{2d}$. Let $H_{\psi}$ be a neural network Hamiltonian with $n$ neurons per layer and $L$ layers, and let $\Phi_{T}$ denote the flow map obtained from the leapfrog integrator with step size $\Delta t$ and $N=T/\Delta t$ steps. Then,
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Phi(z)\|\leq C_{1}\cdot(n\cdot L)^{-1/(2d)}+C_{2}\cdot\Delta t^{2},$
with $C_{1},C_{2}$ constants depending on $\Phi$ and the network architecture.
Proof.
We decompose the overall error into two parts:
$\|\Phi_{T}-\Phi\|\leq\underbrace{\|\Phi_{T}-\Phi_{H}\|}_{\text{Integration error}}+\underbrace{\|\Phi_{H}-\Phi\|}_{\text{Approximation error}},$
where $\Phi_{H}$ denotes the true continuous flow generated by the Hamiltonian $H$ that exactly produces $\Phi$.
(1) Approximation Error:
By the universal approximation theorem for neural networks, a neural network with $n$ neurons per layer and $L$ layers can approximate a smooth function on a compact set with error that decreases as a function of the network size. In our setting, we wish to approximate the underlying Hamiltonian $H$ (or more precisely, its gradient, since the flow is generated by the Hamiltonian vector field $J\nabla H$). Standard results in approximation theory (see, e.g., results on approximation in Sobolev spaces) indicate that, for a function defined on $\mathbb{R}^{2d}$, the error in the $C^{1}$-norm can be bounded by
$\|\nabla H-\nabla H_{\psi}\|_{C^{0}(\Omega)}\leq C_{1}\cdot(n\cdot L)^{-1/(2d)},$
where $C_{1}$ is a constant depending on the smoothness of $H$ and the geometry of $\Omega$. Since the flow $\Phi_{H}$ depends continuously on the vector field, this error propagates to the flow so that
$\|\Phi_{H}-\Phi\|\leq C_{1}\cdot(n\cdot L)^{-1/(2d)}.$
(2) Integration Error:
The leapfrog integrator is a second-order method. This means that, for each integration step, the local truncation error is of order $\mathcal{O}(\Delta t^{3})$, and the global error over $N$ steps accumulates to be of order $\mathcal{O}(\Delta t^{2})$ (since $N\propto 1/\Delta t$ when $T$ is fixed). Therefore, there exists a constant $C_{2}$ (depending on higher derivatives of $H$ and the total integration time $T$) such that
$\|\Phi_{T}-\Phi_{H}\|\leq C_{2}\cdot\Delta t^{2}.$
(3) Combining the Errors:
By the triangle inequality,
$\|\Phi_{T}-\Phi\|\leq\|\Phi_{T}-\Phi_{H}\|+\|\Phi_{H}-\Phi\|\leq C_{2}\cdot\Delta t^{2}+C_{1}\cdot(n\cdot L)^{-1/(2d)}.$
This completes the proof. ∎
6.3 Expressivity Classes
Theorem 6.3 (Expressivity Classes).
Volume-preserving diffeomorphisms on $\mathbb{R}^{2d}$ can be partitioned as follows:
  1. 1.
    Class $\mathcal{C}_{1}$: Exactly representable by flows of quadratic Hamiltonians.
  2. 2.
    Class $\mathcal{C}_{2}$: Efficiently approximable by neural network Hamiltonians of moderate complexity.
  3. 3.
    Class $\mathcal{C}_{3}$: Only approximable by Hamiltonian flows with exponential network complexity.
Moreover, some maps in $\mathcal{C}_{3}$ can be efficiently represented by normalizing flows.
Proof.
We consider each class in turn.
Class $\mathcal{C}_{1}$: Exactly Representable Maps.
Quadratic Hamiltonians have the form
$H(q,p)=\frac{1}{2}z^{T}Az,$
with $z=(q,p)$ and a symmetric matrix $A$. The corresponding Hamiltonian flow is linear and can be written as
$\Phi_{t}(z)=e^{tJA}z,$
where $J$ is the canonical symplectic matrix. Since the exponential of a matrix is computed exactly (or to arbitrary precision) and the mapping is linear, every volume-preserving diffeomorphism that is linear (or that can be exactly represented by such a flow) falls into $\mathcal{C}_{1}$.
Class $\mathcal{C}_{2}$: Efficiently Approximable Maps.
For many smooth volume-preserving diffeomorphisms, the underlying Hamiltonian generating the flow is a smooth function on a compact set. By the universal approximation theorem for neural networks, one can approximate a smooth function to within any $\epsilon>0$ with a neural network whose size grows polynomially in $1/\epsilon$. In particular, there exists a neural network Hamiltonian $H_{\psi}$ with a moderate number of neurons per layer and a moderate number of layers such that
$\sup_{z\in\Omega}\|\nabla H(z)-\nabla H_{\psi}(z)\|<\epsilon.$
Because the flow generated by $H_{\psi}$ depends continuously on the Hamiltonian, the corresponding symplectic flow can approximate the target flow with an error that is also polynomial in the network size. Therefore, maps in $\mathcal{C}_{2}$ are efficiently approximable by neural network Hamiltonians.
Class $\mathcal{C}_{3}$: Hard-to-Approximate Maps.
There exist volume-preserving diffeomorphisms that exhibit highly oscillatory behavior or intricate structures which make the associated Hamiltonian very complex. For such maps, approximating the Hamiltonian $H$ uniformly in the $C^{1}$ norm on a compact set may require a neural network whose size grows exponentially with the dimension $d$ (or with $1/\epsilon$). In these cases, the network complexity is exponential, meaning that these maps can only be approximated by Hamiltonian flows with exponential network complexity. Notably, the structured design of normalizing flows (e.g., using coupling layers that exploit problem structure) can sometimes represent these complex transformations more efficiently than a generic neural network approximation of the Hamiltonian.
Conclusion:
This partitioning illustrates a trade-off between expressivity and computational efficiency. While quadratic Hamiltonians ($\mathcal{C}_{1}$) are exactly representable and many smooth maps ($\mathcal{C}_{2}$) can be efficiently approximated, there exists a class of maps ($\mathcal{C}_{3}$) for which a direct Hamiltonian approximation incurs exponential complexity. Interestingly, normalizing flows may overcome this limitation in certain cases by leveraging structured invertible transformations. ∎
6.4 Extended Universal Approximation for Non-Volume-Preserving Maps
Theorem 6.4 (Extended Universal Approximation).
Let $\Phi:\Omega\to\mathbb{R}^{d}$ be a $C^{1}$-smooth diffeomorphism (not necessarily volume preserving) on a compact set $\Omega$. Then, there exists an SGN-based model that, by incorporating an explicit density correction term, approximates $\Phi$ uniformly to within any $\epsilon>0$.
Proof.
We begin by noting that any $C^{1}$-smooth diffeomorphism $\Phi$ can be decomposed into two components via a factorization:
$\Phi=\Lambda\circ\Psi,$
where:
  • $\Psi:\Omega\to\mathbb{R}^{d}$ is a volume-preserving diffeomorphism, and
  • $\Lambda:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a diffeomorphism that accounts for the non-volume-preserving part of $\Phi$ (often interpreted as a dilation or density adjustment).
Step 1: Approximation of the Volume-Preserving Component.
By Theorem 6.1, for any $\epsilon_{1}>0$ there exists a neural network Hamiltonian $H_{\psi}$ (with Lipschitz continuous gradients) and integration parameters $T>0$ and $\Delta t>0$ (with $N=T/\Delta t$ steps) such that the symplectic flow $\Phi_{T}$ generated by $H_{\psi}$ approximates the volume-preserving map $\Psi$ uniformly on $\Omega$:
$\sup_{z\in\Omega}\|\Phi_{T}(z)-\Psi(z)\|<\epsilon_{1}.$
Step 2: Approximation of the Dilation Component.
Since $\Lambda$ is a $C^{1}$-smooth diffeomorphism on a compact set, standard neural network approximation theorems guarantee that for any $\epsilon_{2}>0$ there exists a feed-forward neural network $\Lambda_{\theta}$ such that
$\sup_{z\in\Psi(\Omega)}\|\Lambda_{\theta}(z)-\Lambda(z)\|<\epsilon_{2}.$
Step 3: Composition and Uniform Approximation.
Define the composed SGN-based model as
$\widetilde{\Phi}(z)=\Lambda_{\theta}\bigl{(}\Phi_{T}(z)\bigr{)}.$
Using the triangle inequality, we have for all $z\in\Omega$:
$\|\widetilde{\Phi}(z)-\Phi(z)\|$ $=\|\Lambda_{\theta}(\Phi_{T}(z))-\Lambda(\Psi(z))\|$
$\leq\|\Lambda_{\theta}(\Phi_{T}(z))-\Lambda_{\theta}(\Psi(z))\|+\|\Lambda_{\theta}(\Psi(z))-\Lambda(\Psi(z))\|$
$\leq L_{\Lambda_{\theta}}\|\Phi_{T}(z)-\Psi(z)\|+\epsilon_{2},$
where $L_{\Lambda_{\theta}}$ is the Lipschitz constant of $\Lambda_{\theta}$. By choosing $\epsilon_{1}$ small enough so that
$L_{\Lambda_{\theta}}\cdot\epsilon_{1}<\epsilon-\epsilon_{2},$
and then selecting $\epsilon_{2}$ such that $\epsilon_{1}+\epsilon_{2}<\epsilon$, we obtain
$\sup_{z\in\Omega}\|\widetilde{\Phi}(z)-\Phi(z)\|<\epsilon.$
Thus, the SGN-based model with the explicit density correction (via the neural network $\Lambda_{\theta}$) uniformly approximates $\Phi$ within any prescribed error $\epsilon>0$. ∎
7 Information-Theoretic Analysis
7.1 Information Geometry of Symplectic Manifolds
Definition 7.1 (Fisher-Rao Metric (Rao 1945)).
For a family $p(x|\theta)$, the Fisher-Rao metric is defined as:
$g_{ij}(\theta)=\mathbb{E}_{p(x|\theta)}\left[\frac{\partial\log p(x|\theta)}{\partial\theta_{i}}\frac{\partial\log p(x|\theta)}{\partial\theta_{j}}\right].$
Figure 4 shows the information geometry on statistical manifolds.
[…]
Figure 4: Information Geometry on Statistical Manifolds. The ellipse represents a statistical manifold endowed with the Fisher-Rao metric, while the two curves illustrate a geodesic and a Hamiltonian flow between two points.
Theorem 7.2 (Symplectic Structure of Exponential Families).
Let $\{p(x|\theta)\}_{\theta\in\Theta}$ be an exponential family. Then, the natural parameter space $\Theta$ admits a symplectic structure with canonical form
$\omega=\sum_{i}d\theta_{i}\wedge d\eta_{i},$
where $\eta_{i}=\frac{\partial\psi(\theta)}{\partial\theta_{i}}$ and $\psi(\theta)$ is the log-partition function. In particular, the second derivatives $\frac{\partial^{2}\psi(\theta)}{\partial\theta_{i}\partial\theta_{j}}$ coincide with the Fisher-Rao metric.
Proof.
An exponential family is expressed as
$p(x|\theta)=h(x)\exp\left(\langle\theta,T(x)\rangle-\psi(\theta)\right),$
where:
  • $\theta\in\Theta$ are the natural parameters,
  • $T(x)$ is the sufficient statistic,
  • $h(x)$ is the base measure, and
  • $\psi(\theta)$ is the log-partition function, defined by
[…]
The mapping $\theta\mapsto\eta$ defined by
$\eta=\nabla\psi(\theta)=\left(\frac{\partial\psi(\theta)}{\partial\theta_{1}},\ldots,\frac{\partial\psi(\theta)}{\partial\theta_{k}}\right)$
is the Legendre transform that sends the natural parameters $\theta$ to the expectation parameters $\eta$.
This Legendre transform is invertible under suitable conditions (which are satisfied in exponential families), thereby establishing a one-to-one correspondence between the coordinates $\theta$ and $\eta$. Consequently, one can introduce a canonical 2-form on the parameter space by
$\omega=\sum_{i}d\theta_{i}\wedge d\eta_{i}.$
We now verify that $\omega$ is a symplectic form, i.e., it is closed and non-degenerate.
Closedness: Since $d\theta_{i}$ and $d\eta_{i}$ are exact 1-forms and the exterior derivative satisfies $d^{2}=0$, we have
$d\omega=d\left(\sum_{i}d\theta_{i}\wedge d\eta_{i}\right)=\sum_{i}d(d\theta_{i}\wedge d\eta_{i})=0.$
Thus, $\omega$ is closed.
Non-degeneracy: In the coordinate system $(\theta_{1},\ldots,\theta_{k},\eta_{1},\ldots,\eta_{k})$, the 2-form
$\omega=\sum_{i=1}^{k}d\theta_{i}\wedge d\eta_{i}$
has full rank $2k$. Hence, for any non-zero vector $v$ in the tangent space, there exists another vector $w$ such that $\omega(v,w)\neq 0$; that is, $\omega$ is non-degenerate.
Next, observe that by differentiating the mapping $\eta_{i}=\frac{\partial\psi(\theta)}{\partial\theta_{i}}$, we obtain
$d\eta_{i}=\sum_{j}\frac{\partial^{2}\psi(\theta)}{\partial\theta_{i}\partial\theta_{j}}d\theta_{j}.$
Thus, in the $\theta$ coordinate chart, the symplectic form can be locally expressed as
$\omega=\sum_{i,j}\frac{\partial^{2}\psi(\theta)}{\partial\theta_{i}\partial\theta_{j}}\,d\theta_{i}\wedge d\theta_{j}.$
The matrix with entries
$g_{ij}(\theta)=\frac{\partial^{2}\psi(\theta)}{\partial\theta_{i}\partial\theta_{j}}$
is known to be positive definite and is, in fact, the Fisher-Rao metric on the parameter space $\Theta$. This connection shows that the canonical symplectic form $\omega$ is intrinsically related to the Fisher-Rao metric.
In summary, the mapping $\theta\mapsto\eta=\nabla\psi(\theta)$ equips the natural parameter space with a symplectic structure given by
$\omega=\sum_{i}d\theta_{i}\wedge d\eta_{i},$
and the Hessian of the log-partition function, $g_{ij}(\theta)$, which defines the Fisher-Rao metric, appears naturally in this context. ∎
Theorem 7.3 (Information-Geometric Interpretation of SGNs).
Assume the latent space of an SGN is endowed with a prior from an exponential family. Then, the Hamiltonian dynamics induced by $H_{\psi}(q,p)$ define geodesic flows on the statistical manifold (with respect to the Fisher-Rao metric).
Proof.
Let the latent space be parameterized by natural parameters $\theta$ of an exponential family. That is, the prior takes the form
$p(x|\theta)=h(x)\exp\left(\langle\theta,T(x)\rangle-\psi(\theta)\right),$
with the log-partition function $\psi(\theta)$ and sufficient statistics $T(x)$. By Theorem 7.2, the natural parameter space $\Theta$ carries a canonical symplectic structure given by
$\omega=\sum_{i}d\theta_{i}\wedge d\eta_{i},$
where $\eta=\nabla\psi(\theta)$. Moreover, the Hessian matrix
$G(\theta)=\nabla^{2}\psi(\theta)$
defines the Fisher-Rao metric on $\Theta$.
Now, consider a Hamiltonian defined on the latent space of the form
$H_{\psi}(q,p)=\frac{1}{2}p^{T}G(q)^{-1}p+V(q),$
where $q$ represents the coordinates corresponding to the natural parameters $\theta$ (or a suitable coordinate representation thereof) and $p$ is the conjugate momentum. Here, $V(q)$ is a potential function that may be chosen to adjust the dynamics; in the simplest case, one may take $V(q)=0$.
The Hamiltonian dynamics are governed by Hamilton’s equations:
$\dot{q}=\frac{\partial H_{\psi}}{\partial p}=G(q)^{-1}p,\quad\dot{p}=-\frac{\partial H_{\psi}}{\partial q}=-\frac{1}{2}p^{T}\frac{\partial\left(G(q)^{-1}\right)}{\partial q}p-\nabla V(q).$
In the special case where $V(q)=0$, the Hamiltonian reduces to a pure kinetic energy term:
$H_{\psi}(q,p)=\frac{1}{2}p^{T}G(q)^{-1}p.$
It is a classical result in Riemannian geometry that the geodesic flow on a manifold with metric $G(q)$ is generated by the Hamiltonian corresponding to the kinetic energy of a free particle (i.e., with no potential term). Therefore, the flow
$\Phi_{t}(q,p)$
generated by this Hamiltonian describes geodesics with respect to the Fisher-Rao metric $G(q)$.
Even if a non-zero potential $V(q)$ is included, for small perturbations the dynamics remain close to geodesic flows or can be interpreted as geodesic flows on a perturbed metric. Hence, the latent dynamics induced by the SGN’s Hamiltonian $H_{\psi}$ correspond to geodesic trajectories on the statistical manifold defined by the prior’s Fisher-Rao metric.
Thus, the Hamiltonian dynamics in SGNs not only provide an invertible and volume-preserving mapping but also offer an intrinsic information-geometric interpretation, as they follow (or approximate) the geodesics of the underlying statistical manifold. ∎
Theorem 7.4 (Information Conservation).
Let $Z_{0}\sim p(z_{0})$ with entropy $H(Z_{0})$ and let $Z_{T}=\Phi_{T}(Z_{0})$ be the transformed latent variable under the symplectic map $\Phi_{T}$. Then:
  1. 1.
    $H(Z_{T})=H(Z_{0})$.
  2. 2.
    For any partition $S\cup S^{c}$ of the coordinates of $Z_{0}$,
    $I(Z_{T}^{S};Z_{T}^{S^{c}})\geq I(Z_{0}^{S};Z_{0}^{S^{c}})-2d\cdot\log(L_{\Phi_{T}}),$
    where $L_{\Phi_{T}}$ is the Lipschitz constant of $\Phi_{T}$.
Proof.
(1) Entropy Preservation:
By the change-of-variables formula for differential entropy, if $Z_{T}=\Phi_{T}(Z_{0})$ is obtained via an invertible mapping $\Phi_{T}$, then
$H(Z_{T})=H(Z_{0})+\mathbb{E}_{Z_{0}}\left[\log\left|\det\frac{\partial\Phi_{T}}{\partial Z_{0}}\right|\right].$
Since $\Phi_{T}$ is symplectic, it preserves volume; that is,
$\left|\det\frac{\partial\Phi_{T}}{\partial Z_{0}}\right|=1\quad\text{for all }Z_{0}.$
Thus, the expectation term vanishes and we have
$H(Z_{T})=H(Z_{0}).$
(2) Mutual Information Bound:
Let $Z_{0}=(Z_{0}^{S},Z_{0}^{S^{c}})$ and $Z_{T}=(Z_{T}^{S},Z_{T}^{S^{c}})$ denote the partition of the latent variable into two complementary subsets of coordinates. The mutual information between the partitions is defined as
$I(Z_{T}^{S};Z_{T}^{S^{c}})=H(Z_{T}^{S})+H(Z_{T}^{S^{c}})-H(Z_{T}).$
Since we have shown $H(Z_{T})=H(Z_{0})$, it suffices to compare the marginal entropies before and after the transformation.
Assume that the mapping $\Phi_{T}$ is Lipschitz continuous with constant $L_{\Phi_{T}}$. Then, standard results in information theory imply that for any Lipschitz mapping $f$ on $\mathbb{R}^{d}$, the change in differential entropy satisfies
$\left|H(f(X))-H(X)\right|\leq d\cdot\log(L_{f}),$
where $d$ is the dimension of the input $X$ and $L_{f}$ is the Lipschitz constant of $f$. Applying this to the marginal transformations of $Z_{0}^{S}$ and $Z_{0}^{S^{c}}$ under $\Phi_{T}$, we obtain
$\left|H(Z_{T}^{S})-H(Z_{0}^{S})\right|\leq d_{S}\cdot\log(L_{\Phi_{T}})\quad\text{and}\quad\left|H(Z_{T}^{S^{c}})-H(Z_{0}^{S^{c}})\right|\leq d_{S^{c}}\cdot\log(L_{\Phi_{T}}),$
where $d_{S}$ and $d_{S^{c}}$ are the dimensions of the partitions $Z_{0}^{S}$ and $Z_{0}^{S^{c}}$, respectively. Since $d_{S}+d_{S^{c}}=2d$, it follows that
$H(Z_{T}^{S})\geq H(Z_{0}^{S})-d_{S}\log(L_{\Phi_{T}})\quad\text{and}\quad H(Z_{T}^{S^{c}})\geq H(Z_{0}^{S^{c}})-d_{S^{c}}\log(L_{\Phi_{T}}).$
Therefore, summing these inequalities gives
$H(Z_{T}^{S})+H(Z_{T}^{S^{c}})\geq H(Z_{0}^{S})+H(Z_{0}^{S^{c}})-(d_{S}+d_{S^{c}})\log(L_{\Phi_{T}}).$
That is,
$H(Z_{T}^{S})+H(Z_{T}^{S^{c}})\geq H(Z_{0}^{S})+H(Z_{0}^{S^{c}})-2d\log(L_{\Phi_{T}}).$
Recall that the mutual information before transformation is
$I(Z_{0}^{S};Z_{0}^{S^{c}})=H(Z_{0}^{S})+H(Z_{0}^{S^{c}})-H(Z_{0}).$
Since $H(Z_{T})=H(Z_{0})$, we have
$I(Z_{T}^{S};Z_{T}^{S^{c}})=H(Z_{T}^{S})+H(Z_{T}^{S^{c}})-H(Z_{T})\geq\left[H(Z_{0}^{S})+H(Z_{0}^{S^{c}})-2d\log(L_{\Phi_{T}})\right]-H(Z_{0}).$
Thus,
$I(Z_{T}^{S};Z_{T}^{S^{c}})\geq I(Z_{0}^{S};Z_{0}^{S^{c}})-2d\log(L_{\Phi_{T}}).$
This completes the proof. ∎
Theorem 7.5 (Optimal Transport Characterization).
Let $p(z_{0})$ and $p(z_{T})$ denote the distributions before and after the symplectic map $\Phi_{T}$. Then, $\Phi_{T}$ is the solution to the optimal transport problem
[…]
with cost
[…]
subject to $z(0)=z_{0}$ and $z(T)=z_{T}$.
Proof.
The proof relies on the dynamic formulation of optimal transport provided by the Benamou–Brenier framework. In this formulation, the optimal transport problem is recast as finding a path $z(t)=(q(t),p(t))$ connecting $z(0)=z_{0}$ to $z(T)=z_{T}$ that minimizes an action integral, subject to the constraint that the time-dependent probability density $p(z,t)$ evolves from $p(z_{0})$ to $p(z_{T})$.
In our setting, the symplectic map $\Phi_{T}$ is generated by Hamiltonian dynamics with Hamiltonian $H_{\psi}(q,p)$. According to Hamilton’s principle, the actual trajectory followed by a system is the one that minimizes (or, more precisely, renders stationary) the action
[…]
Thus, for any pair of endpoints $z_{0}$ and $z_{T}$, the cost to transport $z_{0}$ to $z_{T}$ can be defined as the infimum of this action over all admissible paths:
[…]
where the infimum is taken over all paths $z(\cdot)$ satisfying the boundary conditions $z(0)=z_{0}$ and $z(T)=z_{T}$.
The set $\Pi(p(z_{0}),p(z_{T}))$ consists of all couplings (joint distributions) with marginals $p(z_{0})$ and $p(z_{T})$. The optimal transport problem is then to find a coupling $\gamma$ that minimizes the total cost
[…]
Because the symplectic flow $\Phi_{T}$ precisely transports $p(z_{0})$ to $p(z_{T})$ while following the dynamics that minimize the action (as prescribed by Hamilton’s equations), it follows that $\Phi_{T}$ is the solution to the optimal transport problem with the above cost.
In summary, the symplectic map $\Phi_{T}$ minimizes the action integral
[…]
thereby characterizing it as the optimal transport map between $p(z_{0})$ and $p(z_{T})$ under the cost function $c(z_{0},z_{T})$ defined above. ∎
Theorem 7.6 (Information Bottleneck Optimality).
Consider an SGN with a stochastic encoder $q_{\phi}(z_{0}|x)$ and symplectic flow $\Phi_{T}$. Under suitable conditions on $H_{\psi}$, the model approximates the solution to the information bottleneck problem:
$\min_{p(z|x)}I(X;Z)-\beta I(Z;Y),$
with $Y$ being the target (or reconstructed data) and $\beta$ controlling the trade-off.
Proof.
The SGN is designed with three key components: a stochastic encoder $q_{\phi}(z_{0}|x)$ that maps the input $x$ to an initial latent variable $z_{0}$, a symplectic flow $\Phi_{T}$ that deterministically evolves $z_{0}$ to $z_{T}=\Phi_{T}(z_{0})$, and a decoder that reconstructs or predicts the target $Y$ from the transformed latent variable. A typical training objective is the evidence lower bound (ELBO)
$\mathcal{L}_{\mathrm{SGN}}(x)=\mathbb{E}_{q_{\phi}(z_{0}|x)}\Bigl{[}\log p_{\theta}\bigl{(}x|\Phi_{T}(z_{0})\bigr{)}\Bigr{]}-D_{\mathrm{KL}}\Bigl{(}q_{\phi}(z_{0}|x)\,\|\,p(z_{0})\Bigr{)}.$
Since $\Phi_{T}$ is symplectic, it is invertible and volume preserving; thus, by the change-of-variables formula, the latent representation after the flow, $z_{T}$, satisfies
$I(X;z_{T})=I(X;z_{0}).$
This means that the mutual information between $X$ and the latent representation is fully determined by the encoder $q_{\phi}(z_{0}|x)$.
The KL divergence term $D_{\mathrm{KL}}\bigl{(}q_{\phi}(z_{0}|x)\,\|\,p(z_{0})\bigr{)}$ in the ELBO serves as a regularizer that penalizes the amount of information the latent variable $z_{0}$ carries about $X$. Minimizing this term forces the encoder to compress the representation of $X$, reducing $I(X;z_{0})$. At the same time, the reconstruction (or likelihood) term encourages the preservation of relevant information for predicting $Y$.
This trade-off is precisely what the information bottleneck (IB) principle seeks to balance: it aims to find a representation $Z$ that minimizes $I(X;Z)$ (thus discarding irrelevant information) while preserving as much information as possible about the target $Y$ (maximizing $I(Z;Y)$). The IB objective is typically written as
$\min_{p(z|x)}I(X;Z)-\beta I(Z;Y),$
where $\beta$ is a Lagrange multiplier that controls the trade-off between compression and predictive power.
In the context of SGNs, by appropriately weighting the KL divergence term in the ELBO (or equivalently adjusting hyperparameters such as $\beta$ in an augmented objective), the model is encouraged to learn an encoder that discards non-predictive information while retaining what is necessary to reconstruct $Y$. Due to the invertibility and volume-preserving properties of $\Phi_{T}$, the overall mutual information between $X$ and the latent variable remains preserved after the flow, i.e., $I(X;z_{T})=I(X;z_{0})$.
Therefore, under suitable conditions on the Hamiltonian $H_{\psi}$ (ensuring that the dynamics do not distort the information content) and with an appropriate choice of network architecture and regularization, the SGN approximates the solution to the information bottleneck problem. The model effectively balances the minimization of $I(X;Z)$ (through the KL term) with the maximization of $I(Z;Y)$ (through the reconstruction term), yielding a representation that aligns with the IB objective. ∎
8 Expanded Stability Analysis
8.1 Rigorous Backward Error Analysis
Theorem 8.1 (Modified Hamiltonian).
Let $H_{\psi}(q,p)$ be a $C^{k+1}$-smooth Hamiltonian with $k\geq 3$. Then the leapfrog integrator with step size $\Delta t$ exactly preserves a modified Hamiltonian:
$\tilde{H}(q,p,\Delta t)=H_{\psi}(q,p)+\Delta t^{2}H_{2}(q,p)+\Delta t^{4}H_{4}(q,p)+\ldots+\Delta t^{k-1}H_{k-1}(q,p)+\mathcal{O}(\Delta t^{k+1}).$
Proof.
We begin by expressing the leapfrog integrator as a composition of simpler maps. In particular, the leapfrog update can be written as
$\Phi_{\Delta t}=\exp\left(\frac{\Delta t}{2}L_{H_{\psi}}\right)\circ\exp\left(\Delta t\,L_{H_{\psi}}\right)\circ\exp\left(\frac{\Delta t}{2}L_{H_{\psi}}\right),$
where $L_{H_{\psi}}$ denotes the Lie derivative with respect to the Hamiltonian vector field generated by $H_{\psi}$.
Using the Baker–Campbell–Hausdorff (BCH) formula, we can combine the exponentials in the above composition. The BCH formula tells us that for operators $A$ and $B$,
$\exp(A)\exp(B)=\exp\left(A+B+\frac{1}{2}[A,B]+\frac{1}{12}[A,[A,B]]-\frac{1}{12}[B,[A,B]]+\cdots\right).$
Applying this formula iteratively to the composition in the leapfrog integrator, we obtain an expansion of the form
$\Phi_{\Delta t}=\exp\Bigl{(}\Delta t\,L_{H_{\psi}}+\Delta t^{3}H_{2}+\Delta t^{5}H_{4}+\cdots\Bigr{)},$
where the operators $H_{2},H_{4},\ldots$ are differential operators that depend on higher order commutators of $L_{H_{\psi}}$.
This exponential representation shows that the numerical flow $\Phi_{\Delta t}$ exactly preserves a modified Hamiltonian $\tilde{H}(q,p,\Delta t)$ defined implicitly by the relation
$\exp\left(\Delta t\,L_{\tilde{H}(q,p,\Delta t)}\right)=\Phi_{\Delta t}.$
The expansion of $\tilde{H}$ is given by
$\tilde{H}(q,p,\Delta t)=H_{\psi}(q,p)+\Delta t^{2}H_{2}(q,p)+\Delta t^{4}H_{4}(q,p)+\ldots+\Delta t^{k-1}H_{k-1}(q,p)+\mathcal{O}(\Delta t^{k+1}),$
which shows that the leapfrog integrator exactly preserves this modified Hamiltonian. The accuracy of the expansion, with the remainder term being $\mathcal{O}(\Delta t^{k+1})$, is ensured by the $C^{k+1}$-smoothness of $H_{\psi}$.
Thus, the leapfrog integrator does not exactly preserve the original Hamiltonian $H_{\psi}$, but it exactly preserves the modified Hamiltonian $\tilde{H}(q,p,\Delta t)$ given by the above expansion. ∎
Corollary 8.2 (Energy Conservation).
If $H_{\psi}$ is analytic and the step size $\Delta t$ is sufficiently small, then for exponentially long times $T\leq\exp(c/\Delta t)$,
$\bigl{|}H_{\psi}(q_{T},p_{T})-H_{\psi}(q_{0},p_{0})\bigr{|}\leq C\Delta t^{2},$
with $C$ and $c$ constants independent of $T$ and $\Delta t$.
Proof.
The proof is based on backward error analysis, which shows that a symplectic integrator, such as the leapfrog method, exactly preserves a modified Hamiltonian $\tilde{H}(q,p,\Delta t)$ that can be expanded as
$\tilde{H}(q,p,\Delta t)=H_{\psi}(q,p)+\Delta t^{2}H_{2}(q,p)+\Delta t^{4}H_{4}(q,p)+\ldots+\mathcal{O}(\Delta t^{k+1}),$
where the error expansion holds under the assumption that $H_{\psi}$ is $C^{k+1}$-smooth (and analytic in this corollary).
Because $H_{\psi}$ is analytic, the series converges for sufficiently small $\Delta t$. Thus, $\tilde{H}(q,p,\Delta t)$ remains uniformly close to the original Hamiltonian $H_{\psi}(q,p)$; specifically, the difference
$\left|\tilde{H}(q,p,\Delta t)-H_{\psi}(q,p)\right|$
is bounded by $C_{0}\Delta t^{2}$ for some constant $C_{0}$ independent of $\Delta t$.
Since the leapfrog integrator exactly conserves $\tilde{H}$ along its numerical trajectories, we have
$\tilde{H}(q_{T},p_{T},\Delta t)=\tilde{H}(q_{0},p_{0},\Delta t)$
for the computed states $(q_{T},p_{T})$ and $(q_{0},p_{0})$ at times $T$ and 00, respectively. Therefore,
$\left|H_{\psi}(q_{T},p_{T})-H_{\psi}(q_{0},p_{0})\right|$ $\leq\left|H_{\psi}(q_{T},p_{T})-\tilde{H}(q_{T},p_{T},\Delta t)\right|+\left|\tilde{H}(q_{0},p_{0},\Delta t)-H_{\psi}(q_{0},p_{0})\right|$
$\leq C_{0}\Delta t^{2}+C_{0}\Delta t^{2}=2C_{0}\Delta t^{2}.$
Setting $C=2C_{0}$ establishes the bound.
Moreover, standard results in backward error analysis (see, e.g., Hairer et al.) show that for analytic Hamiltonians, the modified Hamiltonian $\tilde{H}$ is conserved over time intervals that are exponentially long in $1/\Delta t$; that is, for times $T\leq\exp(c/\Delta t)$ for some constant $c>0$.
Thus, the energy drift is bounded by $C\Delta t^{2}$ uniformly for $T\leq\exp(c/\Delta t)$, which demonstrates near energy conservation for sufficiently small $\Delta t$. ∎
8.2 Stability Domains for Neural Network Hamiltonians
Theorem 8.3 (Stability Domains).
Let $H_{\psi}(q,p)$ be a neural network Hamiltonian with Lipschitz continuous gradients satisfying
$\|\nabla_{q}H_{\psi}(q_{1},p_{1})-\nabla_{q}H_{\psi}(q_{2},p_{2})\|\leq L_{q}\|q_{1}-q_{2}\|+L_{qp}\|p_{1}-p_{2}\|,$
$\|\nabla_{p}H_{\psi}(q_{1},p_{1})-\nabla_{p}H_{\psi}(q_{2},p_{2})\|\leq L_{pq}\|q_{1}-q_{2}\|+L_{p}\|p_{1}-p_{2}\|.$
Then, the leapfrog integrator is stable if
$\Delta t<\frac{2}{\sqrt{L_{q}L_{p}+L_{qp}L_{pq}}}.$
Proof.
We analyze the stability of the leapfrog integrator by linearizing its update around a fixed point and then deriving a condition under which the perturbations remain bounded.
The leapfrog scheme for Hamilton’s equations,
$\dot{q}=\nabla_{p}H_{\psi}(q,p),\quad\dot{p}=-\nabla_{q}H_{\psi}(q,p),$
updates the state $(q,p)$ as follows:
$p_{t+\frac{1}{2}}$ $=p_{t}-\frac{\Delta t}{2}\nabla_{q}H_{\psi}(q_{t},p_{t}),$
$q_{t+1}$ $=q_{t}+\Delta t\,\nabla_{p}H_{\psi}\Bigl{(}q_{t},\,p_{t+\frac{1}{2}}\Bigr{)},$
$p_{t+1}$ $=p_{t+\frac{1}{2}}-\frac{\Delta t}{2}\nabla_{q}H_{\psi}\Bigl{(}q_{t+1},\,p_{t+\frac{1}{2}}\Bigr{)}.$
Let $(q^{*},p^{*})$ be a fixed point of the continuous system and define small perturbations
$\delta q_{t}=q_{t}-q^{*},\quad\delta p_{t}=p_{t}-p^{*}.$
Under the Lipschitz assumptions on the gradients of $H_{\psi}$, we have
$\|\nabla_{q}H_{\psi}(q_{t},p_{t})-\nabla_{q}H_{\psi}(q^{*},p^{*})\|\leq L_{q}\|\delta q_{t}\|+L_{qp}\|\delta p_{t}\|,$
$\|\nabla_{p}H_{\psi}(q_{t},p_{t})-\nabla_{p}H_{\psi}(q^{*},p^{*})\|\leq L_{pq}\|\delta q_{t}\|+L_{p}\|\delta p_{t}\|.$
The leapfrog updates can be linearized to obtain a system of the form
$\begin{pmatrix}\delta q_{t+1}\\ \delta p_{t+1}\end{pmatrix}=A\begin{pmatrix}\delta q_{t}\\ \delta p_{t}\end{pmatrix},$
where $A$ is the Jacobian (or update) matrix that depends on the Lipschitz constants and the step size $\Delta t$. Stability of the integrator requires that the spectral radius $\rho(A)$ (the maximum absolute value of the eigenvalues of $A$) satisfies $\rho(A)\leq 1$.
A detailed analysis (see, e.g., Hairer et al.’s work on geometric numerical integration) shows that a sufficient condition for the leapfrog integrator to be stable is that the effective time step $\Delta t$ satisfies
$\Delta t<\frac{2}{\sqrt{L_{q}L_{p}+L_{qp}L_{pq}}}.$
This condition ensures that the eigenvalues of $A$ remain on or within the unit circle, thereby preventing the amplification of errors over iterations.
Thus, under the stated Lipschitz conditions on $\nabla_{q}H_{\psi}$ and $\nabla_{p}H_{\psi}$, the leapfrog integrator is stable provided that
$\Delta t<\frac{2}{\sqrt{L_{q}L_{p}+L_{qp}L_{pq}}}.$
Corollary 8.4 (Neural Network Design for Stability).
If spectral normalization is applied to each weight matrix $W_{l}$ so that $\|W_{l}\|_{2}\leq\sigma$, then the stability condition becomes:
$\Delta t<\frac{2}{\sigma^{L}},$
with $L$ the number of layers.
Proof.
Applying spectral normalization to each weight matrix $W_{l}$ ensures that
$\|W_{l}\|_{2}\leq\sigma\quad\text{for all }l.$
Since the overall Lipschitz constant of a neural network is at most the product of the spectral norms of its layers, it follows that
$L_{\text{net}}\leq\prod_{l=1}^{L}\|W_{l}\|_{2}\leq\sigma^{L}.$
From Theorem 8.3, the leapfrog integrator is stable if
$\Delta t<\frac{2}{\sqrt{L_{q}L_{p}+L_{qp}L_{pq}}}.$
In a neural network Hamiltonian, the Lipschitz constants $L_{q},L_{p},L_{qp},$ and $L_{pq}$ can be collectively bounded by $L_{\text{net}}$ (up to constant factors). Thus, a conservative sufficient condition for stability is
$\Delta t<\frac{2}{L_{\text{net}}}\leq\frac{2}{\sigma^{L}}.$
This condition ensures that the numerical integration via the leapfrog scheme remains stable when using the normalized network, making it a critical design guideline. ∎
8.3 Adaptive Integration Schemes
Theorem 8.5 (Adaptive Step Size with Error Bounds).
Let $\mathcal{E}(q,p,\Delta t)=\|\Phi_{2\Delta t}(q,p)-\Phi_{\Delta t}(\Phi_{\Delta t}(q,p))\|$ be a local error estimator for the leapfrog integrator. If the step size is adapted according to
$\Delta t_{\text{new}}=\Delta t_{\text{old}}\cdot\min\Bigl{(}1.5,\max\bigl{(}0.5,0.9\cdot\Bigl{(}\frac{\tau}{\mathcal{E}}\Bigr{)}^{1/3}\bigr{)}\Bigr{)},$
with target tolerance $\tau$, then:
  1. 1.
    The global error is $\mathcal{O}(\tau)$.
  2. 2.
    The number of steps is asymptotically optimal.
  3. 3.
    Each step preserves the symplectic structure.
Proof.
We analyze the three claims separately.
(1) Global Error is $\mathcal{O}(\tau)$:
For a second-order integrator like the leapfrog method, the local truncation error per step scales as $\mathcal{O}(\Delta t^{3})$. Specifically, the error estimator
$\mathcal{E}(q,p,\Delta t)=\|\Phi_{2\Delta t}(q,p)-\Phi_{\Delta t}(\Phi_{\Delta t}(q,p))\|$
satisfies
$\mathcal{E}(q,p,\Delta t)=K\Delta t^{3},$
for some constant $K$ depending on the higher-order derivatives of the Hamiltonian. The adaptive step size rule
$\Delta t_{\text{new}}=\Delta t_{\text{old}}\cdot\min\Bigl{(}1.5,\max\bigl{(}0.5,0.9\cdot\Bigl{(}\frac{\tau}{\mathcal{E}}\Bigr{)}^{1/3}\bigr{)}\Bigr{)}$
adjusts $\Delta t$ so that the local error $\mathcal{E}$ is approximately equal to the target tolerance $\tau$. Consequently, over the entire integration interval, the cumulative (global) error will be proportional to $\tau$, i.e., $\mathcal{O}(\tau)$.
(2) Asymptotically Optimal Number of Steps:
An adaptive method that adjusts $\Delta t$ to maintain a local error near $\tau$ effectively maximizes the step size subject to the error constraint. This minimizes the total number of steps $N$ required to cover a fixed time interval $T$. Hence, the number of steps is asymptotically optimal in that it is as small as possible while ensuring the local error remains within the target tolerance.
(3) Preservation of the Symplectic Structure:
The leapfrog integrator is symplectic by design; that is, for any step size $\Delta t$, the mapping $\Phi_{\Delta t}$ satisfies
$\left|\det\left(\frac{\partial\Phi_{\Delta t}}{\partial(q,p)}\right)\right|=1.$
Since the adaptive scheme only modifies $\Delta t$ between steps and does not alter the form of the leapfrog update, each step remains symplectic regardless of the chosen step size. Therefore, the overall integration process preserves the symplectic structure exactly at every step.
Conclusion:
The adaptive step size rule ensures that the local error is kept approximately at the target tolerance $\tau$, which in turn guarantees that the global error scales as $\mathcal{O}(\tau)$ and that the number of integration steps is minimized. Moreover, since the leapfrog integrator is inherently symplectic and the adaptation procedure does not alter its structure, each step preserves the symplectic form. This completes the proof. ∎
Theorem 8.6 (Error Bounds by Hamiltonian Class).
Let $\Phi_{T}$ be the flow map computed using the leapfrog integrator with $N$ steps of size $\Delta t=T/N$ applied to a Hamiltonian $H_{\psi}$, and let $\Phi_{H}$ denote the exact flow generated by $H_{\psi}$ over time $T$. Then:
  1. 1.
    For separable Hamiltonians $H_{\psi}(q,p)=K(p)+V(q)$ with $K,V\in C^{3}$,
    $\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{1}\,T\,\Delta t^{2}.$
  2. 2.
    For nearly-integrable Hamiltonians $H_{\psi}(q,p)=H_{0}(q,p)+\epsilon H_{1}(q,p)$ with $\epsilon\ll 1$,
    $\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{2}\,\Bigl{(}T\,\Delta t^{2}+\epsilon\,T\Bigr{)}.$
  3. 3.
    For neural network Hamiltonians with $L$ layers,
    $\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{3}\,L\,T\,\Delta t^{2}.$
Proof.
We decompose the overall error between the numerical flow $\Phi_{T}$ (obtained by the leapfrog integrator) and the exact continuous flow $\Phi_{H}$ into the error incurred at each step, and then sum (or accumulate) these local errors over $N$ steps.
(1) Separable Hamiltonians:
For a separable Hamiltonian of the form
$H_{\psi}(q,p)=K(p)+V(q),$
the leapfrog integrator is a well-known second-order method. That is, the local truncation error per step is of order $\mathcal{O}(\Delta t^{3})$. When this error is accumulated over $N$ steps, the global error grows as $\mathcal{O}(N\Delta t^{3})=\mathcal{O}(T\,\Delta t^{2})$. Therefore, there exists a constant $C_{1}$ (depending on the third derivatives of $K$ and $V$) such that
$\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{1}\,T\,\Delta t^{2}.$
(2) Nearly-Integrable Hamiltonians:
Consider a Hamiltonian of the form
$H_{\psi}(q,p)=H_{0}(q,p)+\epsilon H_{1}(q,p),$
with $\epsilon\ll 1$. In this case, the dominant part $H_{0}$ is integrable and its associated flow can be approximated with a global error of order $\mathcal{O}(T\,\Delta t^{2})$ as in the separable case. The perturbation $\epsilon H_{1}$ introduces an additional error that scales linearly with $\epsilon$ and the evolution time $T$. Thus, the overall error can be bounded by a term $C_{2}\,T\,\Delta t^{2}$ from the integration error plus an extra term $C_{2}\,\epsilon\,T$, leading to
$\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{2}\,\Bigl{(}T\,\Delta t^{2}+\epsilon\,T\Bigr{)}.$
(3) Neural Network Hamiltonians:
When $H_{\psi}$ is represented by a neural network, the integration error not only depends on the step size $\Delta t$ but also on the complexity of the neural network approximator. In particular, if the network has $L$ layers, then the effective Lipschitz constant of the Hamiltonian (and its derivatives) may grow roughly as $\sigma^{L}$ (or more generally, scale linearly with $L$ under appropriate normalization). This increased sensitivity amplifies the local error of the integrator. As a result, the global error becomes
$\mathcal{O}(L\,T\,\Delta t^{2}),$
i.e., there exists a constant $C_{3}$ such that
$\|\Phi_{T}(z_{0})-\Phi_{H}(z_{0})\|\leq C_{3}\,L\,T\,\Delta t^{2}.$
In each case, standard error propagation arguments from numerical analysis of symplectic integrators yield these bounds. The constants $C_{1}$, $C_{2}$, and $C_{3}$ depend on the smoothness of the Hamiltonian $H_{\psi}$ (in particular, on its third derivatives and, in the nearly-integrable case, on the magnitude of the perturbation $\epsilon$), as well as on the specific properties of the neural network architecture in case (3).
This completes the proof. ∎
Theorem 8.7 (Unified Stability Hierarchy).
The stability of SGNs can be characterized at three levels:
  1. 1.
    Integration Stability: The leapfrog integrator is stable with bounded energy error if
    $\Delta t<\frac{2}{\sqrt{L_{H}}},$
    where $L_{H}$ is a Lipschitz constant (or an appropriate measure of the curvature) of the Hamiltonian $H_{\psi}$.
  2. 2.
    Model Stability: The SGN preserves volume and the topological structure of the latent space via its symplectic map. That is, the mapping $\Phi_{T}$ satisfies
    $\left|\det\frac{\partial\Phi_{T}}{\partial z_{0}}\right|=1,$
    ensuring that the transformation does not distort the latent space.
  3. 3.
    Training Stability: Under the condition that the gradients of the objective function (e.g., the ELBO) are Lipschitz continuous with constant $L$, standard convergence results for gradient descent guarantee that, with learning rate
    $\eta<\frac{2}{L},$
    the training procedure converges to a stationary point while preserving the stability properties established in (1) and (2).
Proof.
We prove each level of the stability hierarchy in turn.
(1) Integration Stability:
The leapfrog integrator, being a second-order symplectic method, has local truncation error of order $\mathcal{O}(\Delta t^{3})$ and a global error of order $\mathcal{O}(\Delta t^{2})$. From the error analysis in Theorems 8.1 and 8.5, we know that the integrator remains stable (i.e., the energy error remains bounded) provided that the step size satisfies
$\Delta t<\frac{2}{\sqrt{L_{H}}},$
where $L_{H}$ is a Lipschitz constant associated with the Hamiltonian’s derivatives. This condition ensures that the local linear approximation is non-expansive, thereby keeping the numerical energy close to the true energy over long time intervals.
(2) Model Stability:
By construction, the SGN employs a symplectic integrator to evolve the latent state. The defining property of symplectic maps is that they preserve the canonical symplectic form (and thus volume). Specifically, if $\Phi_{T}$ is the flow map induced by the integrator, then
$\left|\det\frac{\partial\Phi_{T}}{\partial z_{0}}\right|=1.$
This volume preservation implies that the topological structure of the latent space is maintained exactly during the forward and reverse mappings, ensuring model stability.
(3) Training Stability:
The training of SGNs typically involves minimizing an objective function (such as the ELBO) using gradient descent. Under the assumption that the gradients of this objective are Lipschitz continuous with Lipschitz constant $L$, standard results in optimization theory state that gradient descent converges to a stationary point if the learning rate satisfies
$\eta<\frac{2}{L}.$
Thus, with a sufficiently small learning rate, the training process is stable. Moreover, because the training objective is constructed using the symplectic flow $\Phi_{T}$, the favorable stability properties (integration and model stability) are preserved during training.
Conclusion:
Combining these three aspects, we obtain a unified stability hierarchy for SGNs:
  • The integrator is stable if the time step is sufficiently small.
  • The model maintains volume and topological invariance via its symplectic map.
  • The training procedure converges under standard Lipschitz conditions on the gradient, provided an appropriately small learning rate is chosen.
This layered approach to stability ensures that SGNs are robust both in their numerical integration and in their learning dynamics. ∎
9 SGN Training Algorithm
Building on the stability and adaptivity results of Section 8, we present the SGN training procedure. This algorithm optimizes the variational objective (equation 5) by integrating the encoder, symplectic flow, and decoder, while ensuring numerical stability through adaptive step sizes (Theorem 8.5) and the unified stability hierarchy (Theorem 8.7). For neural network Hamiltonians, the stability bound $L_{H}$ may be refined as $L_{q}L_{p}+L_{qp}L_{pq}$ (Theorem 8.3), though we adopt the general form from Theorem 8.7 for simplicity. The procedure is detailed in Algorithm 1.
Algorithm 1 SGN Training Procedure with Adaptive Integration and Stability Guarantees
1:Dataset $\{x^{(i)}\}_{i=1}^{N}$, learning rate $\eta$, total integration time $T$, initial step size $\Delta t_{0}$, tolerance $\tau$, stability bound $L_{H}$, Lipschitz constant of ELBO gradients $L$
2:$\Delta t_{0}<\frac{2}{\sqrt{L_{H}}}$, $\eta<\frac{2}{L}$triangleright Stability conditions (Theorem 8.3), (Theorem 8.7)
3:Initialize parameters $\theta,\phi,\psi$triangleright Decoder, encoder, and Hamiltonian parameters
4:while not converged do
5:     for each minibatch $\{x^{(i)}\}$ do
6:         Compute encoder outputs $(\mu,\Sigma)\leftarrow\text{Encoder}_{\phi}(x)$, where $q_{\phi}(z_{0}|x)=\mathcal{N}(\mu,\Sigma)$
7:         Sample $z_{0}\sim\mathcal{N}(\mu,\Sigma)$triangleright Variational sampling (Section 3.3)
8:         Set $z\leftarrow z_{0}$, $t\leftarrow 0$, $\Delta t\leftarrow\Delta t_{0}$
9:         while $t<T$ do
10:              Split $z$ into canonical coordinates $(q,p)$
11:              $p_{\frac{1}{2}}\leftarrow p-\frac{\Delta t}{2}\nabla_{q}H_{\psi}(q,p)$
12:              $q_{\text{new}}\leftarrow q+\Delta t\,\nabla_{p}H_{\psi}(q,p_{\frac{1}{2}})$
13:              $p_{\text{new}}\leftarrow p_{\frac{1}{2}}-\frac{\Delta t}{2}\nabla_{q}H_{\psi}(q_{\text{new}},p_{\frac{1}{2}})$
14:              Compute local error $\mathcal{E}\leftarrow\bigl{\|}\Phi_{2\Delta t}(q,p)-\Phi_{\Delta t}(\Phi_{\Delta t}(q,p))\bigr{\|}$triangleright Error estimator (Theorem 8.5)
[…]
16:              if $\Delta t\geq\frac{2}{\sqrt{L_{H}}}$ then $\Delta t\leftarrow\frac{1.9}{\sqrt{L_{H}}}$triangleright Integration stability (Theorem 8.3)
17:              Update $z\leftarrow(q_{\text{new}},p_{\text{new}})$, $t\leftarrow t+\Delta t$
18:         end while
19:         Set $z_{T}\leftarrow z$triangleright $z_{T}=\Phi_{T}(z_{0})$, symplectic flow output
20:         Compute reconstruction loss $\ell_{\mathrm{rec}}\leftarrow-\log p_{\theta}(x\mid z_{T})$
21:         Compute KL divergence $\ell_{\mathrm{KL}}\leftarrow D_{\mathrm{KL}}\Bigl{(}q_{\phi}(z_{0}|x)\,\|\,p(z_{0})\Bigr{)}$, where $p(z_{0})=\mathcal{N}(0,I_{2d})$
22:         Accumulate loss $L\leftarrow L+\ell_{\mathrm{rec}}+\ell_{\mathrm{KL}}$triangleright ELBO approximation (equation 5)
23:     end for
24:     Update parameters $(\theta,\phi,\psi)\leftarrow(\theta,\phi,\psi)-\eta\nabla L$triangleright Training stability (Theorem 8.7)
25:end while
10 Conclusion and Future Work
In this work, we have introduced Symplectic Generative Networks (SGNs), a novel framework that integrates Hamiltonian dynamics into deep generative modeling. By leveraging symplectic integrators, SGNs achieve invertibility and exact likelihood evaluation without incurring the high computational overhead associated with Jacobian determinant calculations. Our comprehensive theoretical analysis including strengthened universal approximation results, rigorous information-theoretic foundations, and detailed stability analyses demonstrates that SGNs not only offer computational efficiency relative to normalizing flows but also preserve information more effectively than VAEs.
SGNs offer several key advantages:
  • Exact Likelihood Evaluation: The symplectic structure of the latent-to-data mapping guarantees that the Jacobian determinant is unity, enabling exact likelihood computation.
  • Invertibility and Information Preservation: The volume-preserving property of the symplectic integrator ensures that the transformation is invertible and that the latent representation retains full information from the input.
  • Physical Interpretability: Embedding Hamiltonian dynamics into the model provides a natural interpretation of the evolution of the latent state, making SGNs particularly suitable for applications involving physical conservation laws.
Despite these advantages, several limitations remain. In particular, SGNs exhibit sensitivity to integration parameters and are inherently restricted to representing volume-preserving maps. These constraints may limit their applicability to broader classes of transformations encountered in complex real-world data. Moreover, while our theoretical contributions lay a solid foundation, extensive empirical validation across diverse datasets is still necessary.
Future work will address these challenges by:
  • Extending the SGN framework to accommodate transformations beyond strictly volume-preserving maps, thereby increasing expressivity.
  • Developing adaptive integration strategies and improved network architectures to reduce sensitivity to integration parameters.
  • Conducting comprehensive empirical evaluations to benchmark SGNs against state-of-the-art generative models across various application domains.
Overall, our contributions not only deepen the theoretical understanding of SGNs but also provide practical guidelines for their implementation. We anticipate that future advancements along these lines will further enhance the robustness and applicability of SGNs, establishing them as a competitive approach for deep generative modeling in high-impact, real-world scenarios.